Grade 9Mechanics

a body thrown vertically upwards with initial velocity such that it can reach a maximum height of 15 metre is at the same instance a stone is dropped from a height of 15 metre find the ratio of distances travelled by them when they cross each other

anuj
7 Years agoGrade 9

1 Answer

Arun
7 Years ago
(i) Time of Motion of A = Time of Motion of B = 1/2 Time of Desent of B
 
(ii) Let x be the distnace travelled by the ball
 
(iii) Let y be the distance travelled by the stone
 
(iv) time of Flight = sqrt (2h/g)  where h = x+y
 
(v) Time of desend of stone = qrt (2y/g)
 
(v) from equation III & IV , obtaine relation between x & y
 
(vi) find the ratio of x & y
 
value is 3:1